OpenStudy (anonymous):

What is the equation of the line that passes through the points{1/2,-1/3} and {2,2/3} A. y=2/3x B. y=2/3x-2/3 C. y.3/2x+2/3 D. 1/5x+1/4

OpenStudy (anonymous):

equation of aline is y=mx+c just substitute for the values of x then get Y the equation that satisfies both values is the correct one.

OpenStudy (anonymous):

like for{2,2/3} for B x=2, now y=2x/3-2/3 when you substitute for x you will get y=2/3 hence B correct

OpenStudy (anonymous):

Wait.. Use tow point slope form: \[\frac{y - y_1}{y_2 - y_1} = \frac{x - x_1}{x_2 - x_1}\]

OpenStudy (anonymous):

*two.

OpenStudy (anonymous):

thank guys